# Adversarial Illusions in Multi-Modal Embeddings

### v4

@@ 1 Introduction
Specifically, we show that cross-modal alignment in multi-modal embeddings is highly vulnerable to adversarially generated “illusions.” Given an input $x$, we say that a perturbation $x_{\delta}$ is an illusion if it appears similar to $x$ to a human but aligns with an adversary-chosen $y$ from another modality in the embedding space. An illusion thus “misrepresents” its semantic content to downstream tasks. For example, Figure 2 shows an illusion that aligns an image of Magritte’s famous “This is Not an Apple” painting with the text of Magritte’s quote, “Everything we see hides another thing.” Figures 4 and 4 show other examples. In all cases, perturbed images and audio recordings appear similar to their respective original inputs, but downstream tasks—text and image generation, in these examples—act on the perturbed inputs as if their semantics matched the adversary-chosen texts.
@@
Therefore, the adversary can leverage an attack on a multi-modal embedding into a wholesale compromise of current and future downstream tasks (see Figure 1), including tasks he is unaware of during the attack. We show that multiple tasks—including image generation, text generation, audio retrieval, and different types of zero-shot classification (image, thermal image, and audio)—based on embeddings such as ImageBind, AudioCLIP, OpenCLIP [5], and Amazon’s Titan [^1] https://aws.amazon.com/bedrock/titan/ are all misled by cross-modal illusions. Against zero-shot tasks based on ImageBind and AudioCLIP embeddings, our illusions achieve near-perfect ($>99\%$) attack success rates at standard perturbation bounds.
@@ 2 Background and Related Work
[…]
@@ 3.2 Adversary’s Capabilities
White-box.  Several multi-modal encoders are available as open-sourced code. In this setting, the adversary has access to all of their internal details, including architecture, parameters, and gradients (but not downstream tasks). This level of access allows the adversary to backpropagate gradients from the loss functions defined in Section 4 directly to adversarial illusions, similar to conventional adversarial examples [11, 29].
@@
In Section 4.2 and Section 5.3, we show that backpropagating gradients from multiple surrogates helps generate illusions that succeed against target encoders to which the adversary does not have access.
@@
Most query-based attacks rely on estimating the gradients of the target model, require additional auxiliary information, do not work well for targeted attacks, or fail to scale to a large number of queries [45]. Instead, our attack optimizes over a gradient-free objective (Section 4.3) to create effective adversarial illusions (Section 5.4).
@@
In Section 4.4 and Section 5.4, we show that this strategy, starting with white-box access to surrogate encoders, yields a successful query-only attack on Amazon’s Titan embeddings.
@@ 4 Crafting Cross-Modal Illusions
In the rest of this section, we define the objective functions $\mathcal{L}$ for each threat model from Section 3.
@@ 4.1 White-Box Attack
To optimize Equation 3, we iteratively update perturbation $\delta$ with Projected Gradient Descent (PGD) [29].
@@ 4.3 Query-Based Attack
$\mathcal{L}_{BB}(x_{\delta},y_{t})=-f(x_{\delta},y_{t})-\log\left(\textstyle\sum\nolimits_{y\in\mathcal{Y}nt}e^{f(x_{\delta},y)}\right)$ (5)
@@ 5.1 Setup
  • ImageNet is an image-classification dataset that naturally aligns images and WordNet text labels. Before evaluating on the downstream tasks, each label $y$ is injected into the templates introduced in [33] (i.e., $y\leftarrow\text{``A photo of a \{$y$\}.''}$.”)
@@ 5.2 White-Box Attack Results
Figures 2, 4, 4, 8 and 8 illustrate the effect of our illusions on text and image generation. Figures 4 and 8 show audio illusions against generation and zero-shot classification, respectively. Figures 9 and 10 show thermal-image illusions against zero-shot classification and text generation, respectively. Figure 8 shows how increasing alignment between an audio perturbation and the adversary’s text affects downstream tasks. As cosine similarity between the embeddings increases, the interpretation of the perturbed barking-dog audio changes until it is interpreted as a classical concert by downstream tasks. Figure 8 shows how an image generated from the embedding of an adversarial illusion is classified to the adversary-chosen text by an image classifier.
@@
ImageBind - $100\%$ $0\%$ $0\%$
OpenCLIP-ViT $100\%$ $1\%$ $0\%$
AudioCLIP $0\%$ $0\%$ $92\%$
OpenCLIP-RN50 $0\%$ $0\%$ $88\%$
@@
Zero-shot audio classification and audio retrieval.Tables 4 and 5 demonstrate that our illusions are effective at attacking audio retrieval and classification. For ImageBind, audio data is represented as MEL spectrograms, where decibel levels serve as audio analogs of pixels. Meanwhile, for AudioCLIP, we use time series data of the same structure as in [13]. The optimization procedure is the same as in Section 4.1.
@@ 5.3 Transfer Attack Results
We now show that access to the target encoder’s gradients is not necessary to create adversarial illusions. It may be sufficient to perform a white-box attack on a surrogate encoder and “transfer” the resulting illusions to the target. Table 7 shows that the transfer attack can achieve high zero-shot image classification accuracy, although results depend on the similarity between the surrogate and the target.
Both ImageBind and AudioCLIP add modalities to, respectively, pre-trained ViT and ResNet-50 versions of the popular CLIP embedding, albeit in different ways. In ImageBind, all modalities are mapped into the ViT-based CLIP’s embedding space. AudioCLIP fine-tunes ResNet-50-based CLIP on audiovisual data, thereby “merging” the audio, image, and text embedding spaces. This difference seems to matter. In our experiments, we found that illusions did not transfer to a checkpoint at the end of AudioCLIP’s fine-tuning process but did transfer to a partially trained version provided for compatibility with CLIP (and ostensibly close to CLIP’s embedding space). Table 7 includes this partially trained checkpoint.
With a single surrogate model, we observed that illusions transfer well within but not across model architectures. Even so, as we show in the bottom row of Table 7, this generalizability gap can be addressed by jointly optimizing a single illusion across multiple surrogates (as in Section 4.2). We measured the transfer of illusions generated using both OpenCLIP models as surrogates and found that the same illusions achieve high adversarial alignment in both ImageBind and AudioCLIP’s embedding spaces (see Figure 11). The existence of illusions that attack multiple embedding spaces simultaneously indicates the possibility of universal illusions. We leave this line of research to future work.
@@ 5.4 Query-Based and Hybrid Attack Results
Amazon currently charges \$$\$0.00006$ per image query to the Titan API, so the average cost is \$125.51 per 100 images for the query-based attack and \$108.11 per 100 images for the hybrid attack. For comparison, Google’s Gemini Pro embedding currently charges \$$\$0.0025$ per image query.
@@ 6.1 Feature Distillation
$x$ $x_{\delta}$ $x_{jpeg}$ $x$ $x_{\delta}$ $x_{jpeg}$

### v5

@@ 1 Introduction
Specifically, we show that cross-modal alignment in multi-modal embeddings is highly vulnerable to adversarially generated “illusions.” Given an input $x$, we say that a perturbation $x_{\delta}$ is an illusion if it appears similar to $x$ to a human but aligns with an adversary-chosen $y$ from another modality in the embedding space. An illusion thus “misrepresents” its semantic content to downstream tasks. For example, Figure 2 shows an illusion that aligns an image of Magritte’s famous “This is Not an Apple” painting with the text of Magritte’s quote, “Everything we see hides another thing.” Figures˜4 and 4 show other examples. In all cases, perturbed images and audio recordings appear similar to their respective original inputs, but downstream tasks—text and image generation, in these examples—act on the perturbed inputs as if their semantics matched the adversary-chosen texts.
@@
Therefore, the adversary can leverage an attack on a multi-modal embedding into a wholesale compromise of current and future downstream tasks (see Figure 1), including tasks he is unaware of during the attack. We show that multiple tasks—including image generation, text generation, audio retrieval, and different types of zero-shot classification (image, thermal image, and audio)—based on embeddings such as ImageBind, AudioCLIP, OpenCLIP [5], and Amazon’s Titan [^1] https://aws.amazon.com/bedrock/titan/ are all misled by cross-modal illusions. Against zero-shot tasks based on ImageBind and AudioCLIP embeddings, our illusions achieve near-perfect ($99\%$) attack success rates at standard perturbation bounds.
@@ 2 Background and Related Work
$\text{1.}\;\;\theta(x)$ $=e\;\;\text{ --- embedding of input}$
$\text{2.}\;\;\phi(e)$ $=y\;\;\text{ --- downstream task}$
@@ 3.2 Adversary’s Capabilities
White-box.  Several multi-modal encoders are available as open-sourced code. In this setting, the adversary has access to all of their internal details, including architecture, parameters, and gradients (but not downstream tasks). This level of access allows the adversary to backpropagate gradients from the loss functions defined in Section˜4 directly to adversarial illusions, similar to conventional adversarial examples [11, 29].
@@
In Section˜4.2 and Section˜5.3, we show that backpropagating gradients from multiple surrogates helps generate illusions that succeed against target encoders to which the adversary does not have access.
@@
Most query-based attacks rely on estimating the gradients of the target model, require additional auxiliary information, do not work well for targeted attacks, or fail to scale to a large number of queries [45]. Instead, our attack optimizes over a gradient-free objective (Section˜4.3) to create effective adversarial illusions (Section˜5.4).
@@
In Section˜4.4 and Section˜5.4, we show that this strategy, starting with white-box access to surrogate encoders, yields a successful query-only attack on Amazon’s Titan embeddings.
@@ 4 Crafting Cross-Modal Illusions
In the rest of this section, we define the objective functions $\mathcal{L}$ for each threat model from Section˜3.
@@ 4.1 White-Box Attack
To optimize Equation˜3, we iteratively update perturbation $\delta$ with Projected Gradient Descent (PGD) [29].
@@ 4.3 Query-Based Attack
$\mathcal{L}_{BB}(x_{\delta},y_{t})=-f(x_{\delta},y_{t})+\log\left(\textstyle\sum\nolimits_{y\in\mathcal{Y}nt}e^{f(x_{\delta},y)}\right)$ (5)
@@ 5.1 Setup
  • ImageNet is an image-classification dataset that naturally aligns images and WordNet text labels. Before evaluating on the downstream tasks, each label $y$ is injected into the templates introduced in [33] (i.e., y“A photo of a $y\leftarrow\text${``A photo of a $\{$\$y\$\}.''})
@@ 5.2 White-Box Attack Results
Figures˜2, 4, 4, 8 and 8 illustrate the effect of our illusions on text and image generation. Figures˜4 and 8 show audio illusions against generation and zero-shot classification, respectively. Figures˜9 and 10 show thermal-image illusions against zero-shot classification and text generation, respectively. Figure 8 shows how increasing alignment between an audio perturbation and the adversary’s text affects downstream tasks. As cosine similarity between the embeddings increases, the interpretation of the perturbed barking-dog audio changes until it is interpreted as a classical concert by downstream tasks. Figure 8 shows how an image generated from the embedding of an adversarial illusion is classified to the adversary-chosen text by an image classifier.
@@
ImageBind - $100\%$ $\% 00\%$
OpenCLIP-ViT $0\% - 11\% 00\%$
AudioCLIP $\% 00\%$ - $92\%$
OpenCLIP-RN50 $\% 00\%$ $88\%$ -
@@
Zero-shot audio classification and audio retrieval.Tables˜4 and 5 demonstrate that our illusions are effective at attacking audio retrieval and classification. For ImageBind, audio data is represented as MEL spectrograms, where decibel levels serve as audio analogs of pixels. Meanwhile, for AudioCLIP, we use time series data of the same structure as in [13]. The optimization procedure is the same as in Section˜4.1.
@@ 5.3 Transfer Attack Results
We now show that access to the target encoder’s gradients is not necessary to create adversarial illusions. It may be sufficient to perform a white-box attack on a surrogate encoder and “transfer” the resulting illusions to the target. Table˜7 shows that the transfer attack can achieve high zero-shot image classification accuracy, although results depend on the similarity between the surrogate and the target.
Both ImageBind and AudioCLIP add modalities to, respectively, pre-trained ViT and ResNet-50 versions of the popular CLIP embedding, albeit in different ways. In ImageBind, all modalities are mapped into the ViT-based CLIP’s embedding space. AudioCLIP fine-tunes ResNet-50-based CLIP on audiovisual data, thereby “merging” the audio, image, and text embedding spaces. This difference seems to matter. In our experiments, we found that illusions did not transfer to a checkpoint at the end of AudioCLIP’s fine-tuning process but did transfer to a partially trained version provided for compatibility with CLIP (and ostensibly close to CLIP’s embedding space). Table˜7 includes this partially trained checkpoint.
With a single surrogate model, we observed that illusions transfer well within but not across model architectures. Even so, as we show in the bottom row of Table˜7, this generalizability gap can be addressed by jointly optimizing a single illusion across multiple surrogates (as in Section˜4.2). We measured the transfer of illusions generated using both OpenCLIP models as surrogates and found that the same illusions achieve high adversarial alignment in both ImageBind and AudioCLIP’s embedding spaces (see Figure 11). The existence of illusions that attack multiple embedding spaces simultaneously indicates the possibility of universal illusions. We leave this line of research to future work.
@@ 5.4 Query-Based and Hybrid Attack Results
Amazon currently charges \$0.00006\\$0.00006 per image query to the Titan API, so the average cost is \$125.51 per 100 images for the query-based attack and \$108.11 per 100 images for the hybrid attack. For comparison, Google’s Gemini Pro embedding currently charges \$0.0025\\$0.0025 per image query.
@@ 6.1 Feature Distillation
$x$ $x_{\delta}$ $_{jpeg} xx$ $x_{\delta}$ $x_{jpeg}$